OpenStudy (anonymous):

which function has the range (-infinity, -2] U [0, infinity) explain please!

OpenStudy (anonymous):

@campbell_st please help me again

OpenStudy (anonymous):

a y=cot(2x)-1 b y=sec(x)+1 c y=csc(x)-1 d y=cos(x+1)